Calculate elasticity coefficient for labour demand when wage rate changes from $10 - $15 and quantity of labour changes from 12 units to 6 units


1
Expert's answer
2021-07-05T08:59:52-0400

The elasticity coefficient for labour demand is:

"Ed = \frac{6 - 12}{15 - 10}×\frac{15 + 10}{6 + 12} = -5/3."

So, the demand for labour is elastic.